# What Would an LLM Do? Evaluating Policymaking Capabilities of Large Language Models

### v1

@@ Appendix A Nussbaum’s Central Human Capabilities
We list and describe the 10 central human capabilities in the Capability Approach, based on (Nussbaum 2011).
@@ Appendix B Urban ABM Backbone: Theoretical Foundations
This ABM framework comes from the work of (Aguilera et al. 2024). It relies on the set of Maslow’s need categories ( Physiological, Safety, Belonging, Esteem) as $C$ and the set of needs within each category $c\in C$ as $N_{c}$ (Physiological: food, shelter, sleep, health; Safety: clothing, financial security, employment, education; Belonging: family, friendship, intimacy; Esteem: freedom, status, self-esteem) thus resulting in 14 needs across each category. Each agent $a_{i}\in\mathcal{A}$ maintains:
@@
While informing the behavior of agents by calling LLMs for each decision of each agent during a simulation can be a technical solution instead of relying on a SAT matrix, this approach appears to be constrained by scalability concerns and limited number of tokens context for LLMs. Indeed, numerous LLMs call to simulate the interaction between agents result in a large amount of tokens to be processed and generated by LLMs, which in returns make the simulation costly in terms of money (if you need an access to a private API), energy consumption, and time (Feng et al. 2025).
@@ Appendix C Prompts for LLMs
Prompts for LLM Responses for the Benchmark
@@ 82. Response Structure:
26Context: Barcelona has received 24 million in NextGenerationEU recovery funds earmarked for urgent housing innovation. City planners propose using part of the money to expand APROPBarcelonas experimental shipping-container micro-apartment programso that unsheltered residents near Plaça de les Glòries can move into 92 prefabricated units within eighteen months. Critics argue the funds should instead reinforce permanent Housing First placements or bolster rent-supplement vouchers in adjacent municipalities where rents are lower. Local neighborhood groups worry that another modular block will accelerate gentrification pressures and erase remaining community gardens created on the former elevated ring-road site. With municipal elections a year away, the non-profit Habitat Digne BCN must pick one strategic path to recommend before City Hall freezes the allocation.
@@ Appendix D Additional Experiments
In this section, we include some experimental results that we could not include in the main file due to space constraints.
Comparing LLMs’ Choices
@@
Figure 5: Pairwise comparison of the top choices of various LLMs (with each other as well as a local expert) using heat maps for the three geographic regions with expert assessments. 10 contextualized scenarios are used while comparing experts with LLMs, whereas all 40 contextualized scenarios are considered while comparing LLMs with each other.
We conduct additional experiments regarding location-specific top choices by LLMs on the benchmark. Figure 5 compares top choices by experts and LLMs for each of the three locations assessed by experts – Barcelona, Johannesburg, and South Bend. Only the 10 location-specific contextualized scenarios that were assessed by experts are considered when LLMs are compared with experts, but all 40 contextualized scenarios are considered for comparisons between LLMs. The plots provide additional detail around some of the results outlined in the main file. Note that LLM top choices are more similar to each other and the expert in the locations of Barcelona and South Bend than in Johannesburg.
ABM Policy Comparison
Table 3: Policy Comparison across Scenarios
@@
In this table, we detail the results that are presented in Section 4.3 in terms of needs category for the agents representing PEH per scenario. These results were obtained by modifyng the SAT matrix by calling deepseekr1 API with a temperature of 0.1 and using the prompt presented in C.

### v2

@@ Appendix A Nussbaum’s Central Human Capabilities
We list and describe the 10 central human capabilities in the Capability Approach, based on Nussbaum [2011].
@@ Appendix B Urban ABM Backbone: Theoretical Foundations
This ABM framework comes from the work of Aguilera et al. [2024]. It relies on the set of Maslow’s need categories ( Physiological, Safety, Belonging, Esteem) as $C$ and the set of needs within each category $c\in C$ as $N_{c}$ (Physiological: food, shelter, sleep, health; Safety: clothing, financial security, employment, education; Belonging: family, friendship, intimacy; Esteem: freedom, status, self-esteem) thus resulting in 14 needs across each category. Each agent $a_{i}\in\mathcal{A}$ maintains:
@@
While informing the behavior of agents by calling LLMs for each decision of each agent during a simulation can be a technical solution instead of relying on a SAT matrix, this approach appears to be constrained by scalability concerns and limited number of tokens context for LLMs. Indeed, numerous LLMs call to simulate the interaction between agents result in a large amount of tokens to be processed and generated by LLMs, which in returns make the simulation costly in terms of money (if you need an access to a private API), energy consumption, and time Feng et al. [2025].
@@ Appendix C Prompts for LLMs
C.1 Prompts for LLM Responses for the Benchmark
@@ 82. Response Structure:
26Context: Barcelona has received 24 million in NextGenerationEU recovery funds earmarked for "urgent housing innovation." City planners propose using part of the money to expand APROPBarcelonas experimental shipping-container micro-apartment programso that unsheltered residents near Plaça de les Glòries can move into 92 prefabricated units within eighteen months. Critics argue the funds should instead reinforce permanent Housing First placements or bolster rent-supplement vouchers in adjacent municipalities where rents are lower. Local neighborhood groups worry that another modular block will accelerate gentrification pressures and erase remaining community gardens created on the former elevated ring-road site. With municipal elections a year away, the non-profit Habitat Digne BCN must pick one strategic path to recommend before City Hall freezes the allocation.
@@ Appendix D Additional Experiments
In this section, we include additional experimental results.
D.1 Comparing LLMs’ Choices
@@
Figure 5: Pairwise comparison of the top choices of various LLMs (with each other as well as the local primary expert) using heat maps for the three geographic regions that included expert assessments. 10 contextualized scenarios are used while comparing experts with LLMs, whereas all 40 contextualized scenarios are considered while comparing LLMs with each other.
We conduct additional experiments regarding location-specific top choices by LLMs on the benchmark. Figure 5 compares top choices by experts and LLMs for each of the three locations assessed by experts – Barcelona, Johannesburg, and South Bend. Only the 10 location-specific contextualized scenarios that were assessed by experts are considered when LLMs are compared with experts, but all 40 contextualized scenarios are considered for comparisons between LLMs. The plots provide additional detail around some of the results outlined in the main file. Note that LLM top choices are more similar to each other and the primary expert in the locations of Barcelona and South Bend than in Johannesburg.
D.2 ABM Policy Comparison
Table 3: Comparison of policies across scenarios.
@@ D.2 ABM Policy Comparison
In Table 3, we provide additional detail to the analysis presented in Section 4.3, in terms of needs category for the agents representing PEH per scenario. These results were obtained by modifying the SAT matrix by calling the Deepseekr1 API with a temperature of 0.1 and using the prompt presented in Appendix C.2.